OpenStudy (anonymous):

For what values of c does the curve have maximum and minimum values? F(x) 4x^3+cx^2+4x

OpenStudy (anonymous):

v

OpenStudy (perl):

solve F ' (x) = 0

OpenStudy (anonymous):

@perl how would that work with c? Because it's the anti-discriminant. I have 12x^2+2cx+4=0

OpenStudy (anonymous):

@perl yeah that's what I had so would you do the quadratic, take the value that is greater than 0, and then solve it like a regular min/max by plugging back into the original equation?

OpenStudy (perl):

yes that sounds correct.

OpenStudy (perl):

$$ \Large { f(x) = 4x^3 +cx^2 + 4x \\ f ' (x) = 12x^2 + 2cx + 4 \\ \\ \therefore\\ 12x^2 + 2cx + 4 = 0 \\ \therefore\\ 6x^2 + cx + 2 = 0 \\ \therefore\\ c^2 -4(6)(2)\geq 0 \\ \therefore\\ c^2 \geq 48 } $$